# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ III.8 Sobolev–Lorentz spaces
If $1<p<\infty$, $1<q\leq\infty$, and $1/p+1/p^{\prime}=1$, $1/q+1/q^{\prime}=1$, then the same argument as in [133, Section 1.1.15] implies
@@
Applying estimates for Riesz potentials, we obtain the following embedding results for Sobolev-Lorentz spaces, see for instance [145, Eqs. (1.3)–(1.5)], [13, Ch. 4, Thm. 4.18], and [3, Eq. (1.2.4) and Thm. 3.1.4].
@@
    $\|f\|_{W^{-1,(\frac{n}{n-1},\infty)}(U)}\leq C(n)\|f\|_{L^{1}(U)}.$
@@ IV Weak immersions
Let $\Sigma$ be an $n$-dimensional closed smooth manifold with a fixed smooth reference metric $g_{0}$, $m,k\in\mathbb{N}$, $1\leq p\leq\infty$. We say that $\vec{\Phi}$ is a $W^{k,p}$ weak immersion and write $\vec{\Phi}\in W^{k,p}_{\text{imm}}(\Sigma,\mathbb{R}^{m})$ if $\vec{\Phi}\in W^{k,p}\cap W^{1,\infty}(\Sigma,\mathbb{R}^{m})$ and there exists a constant $\Lambda>0$ such that the following holds for a.e. $q\in\Sigma$ and any $X\in T_{q}\Sigma$:
@@
In this section, we provide some additional material on the notion of weak immersions developed in the lecture notes [168]. First, we prove the existence of conformal coordinates for a weak immersion in Section IV.1. Then, we prove that weak immersions with VMO derivatives have integral densities and are locally injective in Section IV.3. Finally, we prove that a weak immersion can always be approximated by smooth immersions and that the underlying conformal structure (in the case of surfaces) converges as well in Section IV.4. In all this paper, we will denote $D^{2}$ the unit disk of $\mathbb{R}^{2}$.
@@ IV.1 Existence of conformal coordinates
Let $\Sigma$ be a $2$-dimensional closed smooth manifold, $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, $U\subset\Sigma$ be open. Let $\varphi\colon U\subset\Sigma\rightarrow\varphi(U)\subset\mathbb{R}^{2}$ be a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism (i.e. a homeomorphism in $W^{1,2}_{\textup{loc}}$ with a $W^{1,2}_{\textup{loc}}$ inverse). We say $\varphi$ is a (weak) isothermal chart of $\vec{\Phi}$ and $\vec{\Phi}\circ\varphi^{-1}$ is (weakly) conformal if there exists a (Lebesgue) measurable function $\lambda\colon\varphi(U)\rightarrow\mathbb{R}$ such that for $i,j\in\{1,2\}$ there holds
@@
Throughout this paper, we frequently use the notion of non-smooth metrics and their conformal class.
@@
Let $\Sigma$ be an $n$-dimensional smooth manifold. Let $g$ be a measurable section on the bundle $T^{*}\Sigma\otimes T^{*}\Sigma$, we say $g$ is a metric if $g_{p}$ is symmetric and positive definite on $(T_{p}\Sigma)^{2}$ for a.e. $p\in\Sigma$. Let $g_{1},g_{2}$ be two metrics, we say $g_{1}$ is (weakly) conformal to $g_{2}$ if there exists a measurable function $\lambda\colon\Sigma\rightarrow\mathbb{R}$ such that $g_{1}=e^{\lambda}g_{2}$.
@@
    We firstly assume $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ and $\varphi_{1},\varphi_{2}\in W^{1,2}(D^{2},\mathbb{R}^{2})$ be isothermal charts as in Definition IV.2. For $k\in\{1,2\}$, since $\varphi_{k}$ is a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism, by standard degree theory we have $\varphi_{k}$ is differentiable a.e. and either $\det(\nabla\varphi_{k})>0$ a.e. or $\det(\nabla\varphi_{k})<0$ a.e., see [82, Theorem 1.7 & Lemma A.28]. Without loss of generality we assume $\det(\nabla\varphi_{k})>0$ a.e. Let $g_{ij}\coloneqq\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi}$, then following the computation in [90, Section 1.5.1], since $\varphi_{k}$ is an isothermal chart, we have
@@
    The condition $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ implies the right hand side has an $L^{\infty}$ norm strictly less than $1$. In particular, $\varphi_{k}$ is quasiconformal, and $\varphi_{1}\circ(\varphi_{2})^{-1}$ is holomorphic on $\varphi_{2}(D^{2})$, see Definition III.6 and [152, Corollary 1.2.8]. Now return to the case when $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$. Let $\mathcal{V}$ be the set of all the weak isothermal charts. Then by the above argument each transition map between elements of $\mathcal{V}$ is either holomorphic or anti-holomorphic, and $\mathcal{V}$ is an atlas by Theorem IV.4.
@@
    When $\Sigma$ is orientable, there exists a metric $h$ of constant Gaussian curvature $1$, 1-1 or 0 and conformal under $\mathcal{V}$ by the uniformization theorem for compact Riemann surfaces (see e.g. [93, Section 4.4]). When $\Sigma$ is non-orientable, $h$ is constructed by using a $2$-cover from a closed orientable smooth surface to $\Sigma$.
    The fact $\alpha\in W^{2,1}(\Sigma)$ follows from Corollary III.11. As for the embedding $W^{2,1}(\Sigma)\hookrightarrow C^{0}(\Sigma)$, see for instance [80, Theorem 3.3.4 & 3.3.10].
@@
Using the conformal structure associated to $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, the relation (I.6) follows from standard computations, see for instance [168, Theorem 1.6]. By integrating (I.6) with respect to $d\textup{vol}_{h}$ (when $\Sigma$ is orientable), we have (I.7) by the Gauss–Bonnet theorem for the smooth metric $h$. Thus Theorem I.1 is a consequence of Corollary IV.5.
The proof of Theorem IV.4 is similar to the case when $\vec{\Phi}$ is a smooth immersion, while the major difficulty lies in the following analytical lemma, which is a weak inverse function theorem, and its higher-dimensional generalization is proved in [78]. It does not hold if we replace $W^{2,2}$ by some lower regularity from the perspective of Sobolev spaces. One counter example is the map $z\mapsto z^{2}/|z|$. Here we provide a different proof for the $2$-dimensional case. Instead of using properties of BLD maps [130, Lemma 4.6 & Theorem 4.7], we use regularity results for Beltrami equations [42]. Thanks to the embedding $W^{1,2}(D^{2})\hookrightarrow VMO(D^{2})$, we obtain another proof in Corollary IV.13.
Lemma IV.7 ([78, Theorem 1.1]).
@@
Let $h\coloneqq\phi\circ f^{-1}$. By the same computation as in [152, Corollary 1.2.8], the map $h$ is holomorphic on $f(D_{1/2}(0))$. Let $k\geq 1$ be the order of zero of $h$ at 0. Then there exists a constant $A\neq 0\in\mathbb{C}$, $0\in U_{1}\subset f(D_{1/2}(0))$, a map $g\in C^{\infty}(\overline{U_{1}},\mathbb{C})$ conformal from the bounded open set $U_{1}$ to $g(U_{1})\subset\mathbb{C}$ with $g(0)=0$ and $g^{\prime}(0)=1$ such that $h=A\,g^{k}$ on $U_{1}$. Let $f_{0}\coloneqq g\circ f$, then since $h=\phi\circ f^{-1}$, we have $\phi=h\circ f=A(g\circ f)^{k}=A\,f_{0}^{k}$ on $f^{-1}(U_{1})$. Since $f$ is quasiconformal and satisfies the equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$ with $\mu\in W^{1,2}(\mathbb{C})$ of compact support, by Lemma III.8 we have $f\in W^{2,q}_{\textup{loc}}(\mathbb{C})$ for any $q<2$. Then by Sobolev embeddings we have $f\in C^{0,\alpha}_{\textup{loc}}(\mathbb{C})$ for any $\alpha<1$. In particular, $f_{0}=g\circ f\in C^{0,\alpha}(f^{-1}(\overline{U_{1}}))$, and it holds that
$|f_{0}(x)|\leq C(\alpha)|x|^{\alpha},\qquad\text{for all }x\in f^{-1}(\overline{U_{1}})\,\text{ and all }\alpha<1.$ (IV.3)
Now we prove $k=1$. Suppose $k\geq 2$, then there exists a positive constant $C_{1}$ such that
@@
Fix $\alpha>\frac{1}{2}$. By (IV.3) and (IV.4), there exists a positive constant $C_{2}$ such that for any $x\in f^{-1}(\overline{U_{1}})$, it holds
@@
Since $4\alpha>2$, (IV.5) implies
$\int_{f^{-1}(U_{1})}|\partial_{z}f_{0}|^{4}=\infty.$
But by Sobolev embedding again, we have $f\in W^{1,p}_{\textup{loc}}(\mathbb{C})$ for any $p<\infty$, hence we obtain $f_{0}=g\circ f\in W^{1,4}(f^{-1}(U_{1}))$, a contradiction arises.
Let $0\in U\Subset f^{-1}(U_{1})$. Since $k=1$, we have $\phi=A(g\circ f)$ is quasiconformal on $U$, and $\phi^{-1}\in W^{1,\infty}\cap W^{2,2}(\phi(U),\mathbb{R}^{2})$ since $D\phi^{-1}=(D\phi)^{-1}\circ\phi^{-1}$ a.e. on $\phi(U)$. ∎
We are now ready to prove Theorem IV.4
@@
$\vec{f}_{1}\coloneqq\frac{\partial_{1}\vec{\Phi}}{|\partial_{1}\vec{\Phi}|},\qquad\text{ and }\qquad\vec{f}_{2}\coloneqq\frac{\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}}{|\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}|}.$ (IV.6)
@@
$\vec{f}^{\,\theta}_{1}+i\vec{f}^{\,\theta}_{2}=e^{i\theta}(\vec{f}_{1}+i\vec{f}_{2}),$
then we still have $(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}(\mathbb{R}^{m}))$. We denote $g=g_{\vec{\Phi}}$. We look now for a rotation of this orthonormal frame realizing the following absolute minimum (for notation, see xvii)
$\begin{array}[]{l}\displaystyle\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1},d\vec{f}^{\,\theta}_{2}\rangle|^{2}_{g}\ d\textup{vol}_{g}\\[14.22636pt] \displaystyle=\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}\,\langle\vec{f}^{\,\theta}_{1},\partial_{i}\vec{f}^{\,\theta}_{2}\rangle\,\langle\vec{f}^{\,\theta}_{1},\partial_{j}\vec{f}^{\,\theta}_{2}\rangle\,\sqrt{\det g}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.8)
@@
$\langle\vec{f}^{\,\theta}_{1},\partial_{k}\vec{f}^{\,\theta}_{2}\rangle=\partial_{k}\theta+\langle\vec{f}_{1},\partial_{k}\vec{f}_{2}\rangle.$ (IV.9)
@@
$\int_{D^{2}}d\phi\wedge*_{g}\left(d\theta+\langle\vec{f}_{1},d\vec{f}_{2}\rangle\right)=0,\qquad\text{for all }\,\phi\in W^{1,2}(D^{2},{\mathbb{R}}).$ (IV.10)
Denote $(\vec{e}_{1},\vec{e}_{2})\coloneqq(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})$ for this value of $\theta$. From (IV.10) we obtain that
@@
$d\lambda=*_{g}\langle\vec{e}_{1},d\vec{e}_{2}\rangle.$ (IV.11)
Substituting this into (IV.10), we have
$\int_{D^{2}}d\phi\wedge d\lambda=0,\qquad\text{for all }\,\phi\in W^{1,2}(D^{2},{\mathbb{R}}).$
@@
It follows that $\lambda\big|_{\mathbb{S}^{1}}$ is constant by fundamental lemma of the calculus of variations [95, Lemma 1.1.1]. Up to adding a constant we can assume $\lambda\in W^{1,2}_{0}(D^{2})$. By applying the operator $_{g}\,d\,*_{g}$ to the equation (IV.11), we obtain
@@
We deduce from Lemma III.3 that $\lambda\in C^{0}(\overline{D^{2}})\cap W^{1,2}_{0}(D^{2},{\mathbb{R}})$. For $i\in\{1,2\}$, we define pullback of $\vec{e}_{i}$ by $\vec{\Phi}$ and the dual $1$-forms
$\begin{dcases}e_{i}\coloneqq g^{jk}\,\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,\dfrac{\partial}{\partial x^{k}}\in L^{\infty}\cap W^{1,2}(D^{2},TD^{2}),\\[5.69054pt] e_{i}^{\ast}\coloneqq\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,dx^{j}=\langle\vec{e}_{i},d\vec{\Phi}\rangle\in L^{\infty}\cap W^{1,2}(D^{2},T^{*}D^{2}).\end{dcases}$ (IV.12)
It holds that $e_{i}^{*}(e_{j})=\delta_{ij}$ and $d\vec{\Phi}(e_{i})=\vec{e}_{i}$ for any $i,j\in\{1,2\}$. Since both $e_{i}^{*}$ and $e_{i}$ are in $L^{\infty}\cap W^{1,2}(D^{2})$, by a weak version of Cartan’s magic formula, we have as in the smooth case [168, Equation (1.70)]:
$de_{i}^{*}=-e_{i}^{*}([e_{1},e_{2}])\,e_{1}^{*}\wedge e_{2}^{*}=d\lambda\wedge e_{i}^{\ast}.$ (IV.13)
@@
where $\pi_{T}$ is the orthogonal projection onto $d\vec{\Phi}(TD^{2})$. Since $\lambda\in L^{\infty}(D^{2})\cap W^{1,2}(D^{2},{\mathbb{R}})$, we deduce from the chain rule that $e^{-\lambda}\in W^{1,2}(D^{2})$ and $d(e^{-\lambda})=-e^{-\lambda}\,d\lambda\in L^{2}(D^{2},T^{*}D^{2})$. Hence the product rule and (IV.13) lead to
@@
$d\varphi^{i}=e^{-\lambda}\,e_{i}^{*}.$ (IV.14)
Let $\varphi\coloneqq(\varphi^{1},\varphi^{2})\in W^{1,\infty}\cap W^{2,2}(D^{2},{\mathbb{R}^{2}})$. By (IV.12) and (IV.14), we have
@@
Since $(e_{1},e_{2})$ is positive and using (IV.7), we obtain that all the eigenvalues of $(\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi})_{1\leq i,j\leq 2}$ lie in $\Lambda^{-1},\Lambda]$ and thus, it holds that
@@
IV.2 Constant of Wente inequality for a metric and consequence

### v3

@@ III.8 Sobolev–Lorentz spaces
If $1<p<\infty$, $1<q\leq\infty$, and $1/p+1/p^{\prime}=1$, $1/q+1/q^{\prime}=1$, then the same argument as in [137, Section 1.1.15] implies
@@
Applying estimates for Riesz potentials, we obtain the following embedding results for Sobolev–Lorentz spaces, see for instance [148, Eqs. (1.3)–(1.5)], [13, Ch. 4, Thm. 4.18], and [3, Eq. (1.2.4) and Thm. 3.1.4].
@@
    $\|f\|_{W^{-1,\left(\frac{n}{n-1},\infty\right)}(U)}\leq C(n)\|f\|_{L^{1}(U)}.$
@@ IV Weak immersions
Let $\Sigma$ be an $n$-dimensional closed smooth manifold with a fixed smooth reference metric $g_{0}$. Fix $m,k\in\mathbb{N}$ and $1\leq p\leq\infty$. We say that $\vec{\Phi}$ is a $W^{k,p}$ weak immersion and write $\vec{\Phi}\in W^{k,p}_{\text{imm}}(\Sigma,\mathbb{R}^{m})$ if $\vec{\Phi}\in W^{k,p}\cap W^{1,\infty}(\Sigma,\mathbb{R}^{m})$ and there exists a constant $\Lambda>0$ such that the following holds for a.e. $q\in\Sigma$ and any $X\in T_{q}\Sigma$:
@@
In this section, we provide some additional material on the notion of weak immersions developed in the lecture notes [171]. First, we prove the existence of conformal coordinates for a weak immersion in Section IV.1. Then, we prove that weak immersions with VMO derivatives have integral densities and are locally injective in Section IV.3. Finally, we prove that a weak immersion can always be approximated by smooth immersions and that the underlying conformal structure (in the case of surfaces) converges as well in Section IV.4. Throughout this paper, we denote by $D^{2}$ the unit disk of $\mathbb{R}^{2}$.
@@ IV.1 Existence of conformal coordinates
Let $\Sigma$ be a $2$-dimensional closed smooth manifold. Consider a weak immersion $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$ and $U\subset\Sigma$ an open set. Let $\varphi\colon U\subset\Sigma\rightarrow\varphi(U)\subset\mathbb{R}^{2}$ be a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism (i.e. a homeomorphism in $W^{1,2}_{\textup{loc}}$ with a $W^{1,2}_{\textup{loc}}$ inverse). We say $\varphi$ is a (weak) isothermal chart of $\vec{\Phi}$ and $\vec{\Phi}\circ\varphi^{-1}$ is (weakly) conformal if there exists a (Lebesgue) measurable function $\lambda\colon\varphi(U)\rightarrow\mathbb{R}$ such that for $i,j\in\{1,2\}$ there holds
@@
Throughout this paper, we will frequently use the notion of non-smooth metrics and their conformal class.
@@
Let $\Sigma$ be an $n$-dimensional smooth manifold. Let $g$ be a measurable section on the bundle $T^{*}\Sigma\otimes T^{*}\Sigma$, we say that $g$ is a metric if $g_{p}$ is symmetric and positive definite on $(T_{p}\Sigma)^{2}$ for a.e. $p\in\Sigma$. Let $g_{1},g_{2}$ be two metrics, we say that $g_{1}$ is (weakly) conformal to $g_{2}$ if there exists a measurable function $\lambda\colon\Sigma\rightarrow\mathbb{R}$ such that $g_{1}=e^{\lambda}g_{2}$.
@@
    We first assume $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ and $\varphi_{1},\varphi_{2}\in W^{1,2}(D^{2},\mathbb{R}^{2})$ be isothermal charts as in Definition IV.2. For $k\in\{1,2\}$, since $\varphi_{k}$ is a $W^{1,2}_{\textup{loc}}$ bi-Sobolev homeomorphism, by standard degree theory we have $\varphi_{k}$ is differentiable a.e. and either $\det(\nabla\varphi_{k})>0$ a.e. or $\det(\nabla\varphi_{k})<0$ a.e., see [83, Theorem 1.7 & Lemma A.28]. Without loss of generality we assume $\det(\nabla\varphi_{k})>0$ a.e. Let $g_{ij}\coloneqq\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi}$, then following the computation in [91, Section 1.5.1], since $\varphi_{k}$ is an isothermal chart, we have
@@
    The condition $\vec{\Phi}\in W^{1,\infty}_{\textup{imm}}(D^{2},\mathbb{R}^{m})$ implies the right-hand side has an $L^{\infty}$ norm strictly less than $1$. In particular, $\varphi_{k}$ is quasiconformal, and $\varphi_{1}\circ(\varphi_{2})^{-1}$ is holomorphic on $\varphi_{2}(D^{2})$, see Definition III.6 and [155, Corollary 1.2.8]. Now return to the case when $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$. Let $\mathcal{V}$ be the set of all the weak isothermal charts. Then by the above argument each transition map between elements of $\mathcal{V}$ is either holomorphic or anti-holomorphic, and $\mathcal{V}$ is an atlas by Theorem IV.4.
@@
    When $\Sigma$ is orientable, there exists a metric $h$ of constant Gaussian curvature $1$, 1-1 or 0 and conformal under $\mathcal{V}$ by the uniformization theorem for compact Riemann surfaces (see e.g. [94, Section 4.4]). When $\Sigma$ is non-orientable, $h$ is constructed by using a $2$-cover from a closed orientable smooth surface to $\Sigma$.
    The fact $\alpha\in W^{2,1}(\Sigma)$ follows from Corollary III.11. Concerning the embedding $W^{2,1}(\Sigma)\hookrightarrow C^{0}(\Sigma)$ we refer to [81, Theorem 3.3.4 & 3.3.10].
@@
Using the conformal structure associated to $\vec{\Phi}\in W^{2,2}_{\textup{imm}}(\Sigma,\mathbb{R}^{m})$, the relation (I.6) follows from standard computations, see for instance [171, Theorem 1.6]. By integrating (I.6) with respect to $d\textup{vol}_{h}$ (when $\Sigma$ is orientable), we have (I.7) by the Gauss–Bonnet theorem for the smooth metric $h$. Thus Theorem I.1 is a consequence of Corollary IV.5.
The proof of Theorem IV.4 is similar to the case when $\vec{\Phi}$ is a smooth immersion, while the major difficulty lies in the following analytical lemma, which is a weak inverse function theorem, and its higher-dimensional generalization is proved in [79]. It does not hold if we replace $W^{2,2}$ by some lower regularity from the perspective of Sobolev spaces. One counter example is the map $z\mapsto z^{2}/|z|$. Here we provide a different proof for the $2$-dimensional case. Instead of using properties of maps with bounded length distortion (BLD maps) [134, Lemma 4.6 & Theorem 4.7], we use regularity results for Beltrami equations [42]. Thanks to the embedding $W^{1,2}(D^{2})\hookrightarrow VMO(D^{2})$, we obtain another proof in Corollary IV.13.
Lemma IV.7 ([79, Theorem 1.1]).
@@
Let $h\coloneqq\phi\circ f^{-1}$. By the same computation as in [155, Corollary 1.2.8], the map $h$ is holomorphic on $f(D_{1/2}(0))$. Let $k\geq 1$ be the order of zero of $h$ at 0. Then there exists a constant $A\neq 0\in\mathbb{C}$, a bounded open set $U_{1}\subset f(D_{1/2}(0))$ containing 0 and a map $g\in C^{\infty}(\overline{U_{1}},\mathbb{C})$ conformal from $U_{1}$ to $g(U_{1})\subset\mathbb{C}$ with $g(0)=0$ and $g^{\prime}(0)=1$ such that $h=A\,g^{k}$ on $U_{1}$. Let $f_{0}\coloneqq g\circ f$, then since $h=\phi\circ f^{-1}$, we have $\phi=h\circ f=A(g\circ f)^{k}=A\,f_{0}^{k}$ on $f^{-1}(U_{1})$. Since $f$ is quasiconformal and satisfies the equation $\partial_{\bar{z}}f=\mu\,\partial_{z}f$ with $\mu\in W^{1,2}(\mathbb{C})$ of compact support, by Lemma III.8 we have $f\in W^{2,q}_{\textup{loc}}(\mathbb{C})$ for any $q<2$. Fix $\alpha\in(0,1)$. By the Sobolev embedding we have $f\in C^{0,\alpha}_{\textup{loc}}(\mathbb{C})$. In particular, this implies that $f_{0}=g\circ f\in C^{0,\alpha}(f^{-1}(\overline{U_{1}}))$, and there exists a constant $C_{0}>0$ such that for all $x\in f^{-1}\big(\overline{U_{1}}\big)$, it holds that
$|f_{0}(x)|\leq C_{0}|x|^{\alpha}.$ (IV.3)
Now we prove that $k=1$. By contradiction, suppose that $k\geq 2$. Then there exists a positive constant $C_{1}$ such that
@@
By (IV.3) and (IV.4), there exists a positive constant $C_{2}$ such that for all $x\in f^{-1}(\overline{U_{1}})$, it holds that
@@
Choose $p\in(1,\infty)$ such that $\alpha p\geq 2$. Then it follows from (IV.5) that
$\int_{f^{-1}(U_{1})}|\partial_{z}f_{0}|^{p}=\infty.$ (IV.6)
Since $f\in W^{2,q}_{\textup{loc}}(\mathbb{C})$ for any $q<2$, we obtain by Sobolev embedding that $f\in W^{1,p}_{\textup{loc}}(\mathbb{C})$. Hence we have $f_{0}=g\circ f\in W^{1,p}(f^{-1}(U_{1}))$, which contradicts (IV.6).
Consider an open set $U$ such that $0\in U\Subset f^{-1}(U_{1})$. Since $k=1$, we have that $\phi=A(g\circ f)$. Hence, $\phi$ is quasiconformal on $U$, and $\phi^{-1}\in W^{1,\infty}\cap W^{2,2}(\phi(U),\mathbb{R}^{2})$ since $D\phi^{-1}=(D\phi)^{-1}\circ\phi^{-1}$ a.e. on $\phi(U)$. ∎
We are now ready to prove Theorem IV.4.
@@
$\vec{f}_{1}\coloneqq\frac{\partial_{1}\vec{\Phi}}{|\partial_{1}\vec{\Phi}|},\qquad\text{ and }\qquad\vec{f}_{2}\coloneqq\frac{\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}}{|\partial_{2}\vec{\Phi}-(\partial_{2}\vec{\Phi}\cdot\vec{f}_{1})\vec{f}_{1}|}.$ (IV.7)
@@
$\vec{f}^{\,\theta}_{1}+i\vec{f}^{\,\theta}_{2}=e^{i\theta}(\vec{f}_{1}+i\vec{f}_{2}).$
We have $(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})\in L^{\infty}\cap W^{1,2}(D^{2},V_{2}(\mathbb{R}^{m}))$. We denote $g=g_{\vec{\Phi}}$. We look now for a rotation of this orthonormal frame realizing the following absolute minimum (for notation, see xvii)
$\begin{array}[]{l}\displaystyle\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}|\langle\vec{f}^{\,\theta}_{1},d\vec{f}^{\,\theta}_{2}\rangle|^{2}_{g}\ d\textup{vol}_{g}\\[14.22636pt] \displaystyle=\inf_{\theta\in W^{1,2}(D^{2},{\mathbb{R}})}\int_{D^{2}}\sum_{i,j=1}^{2}g^{ij}\,\langle\vec{f}^{\,\theta}_{1},\partial_{i}\vec{f}^{\,\theta}_{2}\rangle\,\langle\vec{f}^{\,\theta}_{1},\partial_{j}\vec{f}^{\,\theta}_{2}\rangle\,\sqrt{\det g}\ dx^{1}\wedge dx^{2}.\end{array}$ (IV.9)
@@
$\langle\vec{f}^{\,\theta}_{1},\partial_{k}\vec{f}^{\,\theta}_{2}\rangle=\partial_{k}\theta+\langle\vec{f}_{1},\partial_{k}\vec{f}_{2}\rangle.$ (IV.10)
@@
$\forall\phi\in W^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge*_{g}\left(d\theta+\langle\vec{f}_{1},d\vec{f}_{2}\rangle\right)=0.$ (IV.11)
Denote $(\vec{e}_{1},\vec{e}_{2})\coloneqq(\vec{f}^{\,\theta}_{1},\vec{f}^{\,\theta}_{2})$ for this value of $\theta$. From (IV.11) we obtain that
@@
$d\lambda=*_{g}\langle\vec{e}_{1},d\vec{e}_{2}\rangle.$ (IV.12)
Substituting this into (IV.11), we have
$\forall\phi\in W^{1,2}(D^{2},{\mathbb{R}}),\qquad\int_{D^{2}}d\phi\wedge d\lambda=0.$
@@
It follows that $\lambda\big|_{\mathbb{S}^{1}}$ is constant by the fundamental lemma of the calculus of variations [96, Lemma 1.1.1]. Up to adding a constant we can assume $\lambda\in W^{1,2}_{0}(D^{2})$. By applying the operator $_{g}\,d\,*_{g}$ to the equation (IV.12), we obtain
@@
We deduce from Lemma III.3 that $\lambda\in C^{0}(\overline{D^{2}})\cap W^{1,2}_{0}(D^{2},{\mathbb{R}})$. For $i\in\{1,2\}$, we define the pullback of $\vec{e}_{i}$ by $\vec{\Phi}$ and the dual $1$-forms
$\begin{dcases}e_{i}\coloneqq g^{jk}\,\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,\dfrac{\partial}{\partial x^{k}}\in L^{\infty}\cap W^{1,2}(D^{2},TD^{2}),\\[5.69054pt] e_{i}^{\ast}\coloneqq\langle\vec{e}_{i},\partial_{j}\vec{\Phi}\rangle\,dx^{j}=\langle\vec{e}_{i},d\vec{\Phi}\rangle\in L^{\infty}\cap W^{1,2}(D^{2},T^{*}D^{2}).\end{dcases}$ (IV.13)
It holds that $e_{i}^{*}(e_{j})=\delta_{ij}$ and $d\vec{\Phi}(e_{i})=\vec{e}_{i}$ for any $i,j\in\{1,2\}$. Since both $e_{i}^{*}$ and $e_{i}$ are in $L^{\infty}\cap W^{1,2}(D^{2})$, by a weak version of Cartan’s magic formula, we have as in the smooth case [171, Equation (1.70)]:
$de_{i}^{*}=-e_{i}^{*}([e_{1},e_{2}])\,e_{1}^{*}\wedge e_{2}^{*}=d\lambda\wedge e_{i}^{\ast}.$ (IV.14)
@@
where $\pi_{T}$ is the orthogonal projection onto $d\vec{\Phi}(TD^{2})$. Since $\lambda\in L^{\infty}(D^{2})\cap W^{1,2}(D^{2},{\mathbb{R}})$, we deduce from the chain rule that $e^{-\lambda}\in W^{1,2}(D^{2})$ and $d(e^{-\lambda})=-e^{-\lambda}\,d\lambda\in L^{2}(D^{2},T^{*}D^{2})$. Hence the product rule and (IV.14) lead to
@@
$d\varphi^{i}=e^{-\lambda}\,e_{i}^{*}.$ (IV.15)
Let $\varphi\coloneqq(\varphi^{1},\varphi^{2})\in W^{1,\infty}\cap W^{2,2}(D^{2},{\mathbb{R}^{2}})$. By (IV.13) and (IV.15), we have
@@
Since $(e_{1},e_{2})$ is positive and using (IV.8), we obtain that all the eigenvalues of $(\partial_{i}\vec{\Phi}\cdot\partial_{j}\vec{\Phi})_{1\leq i,j\leq 2}$ lie in $\Lambda^{-1},\Lambda]$ and thus, it holds that
@@
IV.2 Constant of Wente’s inequality for a metric and consequences